Novak Corp. has 10,100 shares of 9%, $100 par value, cumulative preferred stock outstanding at December 31, 2022. No dividends were declared in 2020 or 2021. If Novak wants to pay $395,000 of dividends in 2022, what amount of dividends will common stockholders receive?

Dividends allocated to common stock
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Novak
Calculation of cumulative Preference dividends Amount $ Note
Number of Preference shares         10,100.00 A
Par Value               100.00 B
Value of Preference shares 1,010,000.00 C=A*B
Divided rate 9% D
Annual Divided amount         90,900.00 E=C*D
Preference Divided for 2020         90,900.00 See E
Preference Divided for 2021         90,900.00 See E
Preference Divided for 2022         90,900.00 See E
Cumulative Preference dividends       272,700.00 F
Total dividends payable in 2022       395,000.00 G
Less: Cumulative Preference dividends for 3 years from 2020 to 2022       272,700.00 See F
Common stock dividend       122,300.00 H=G-F
So Common stockholders will receive $ 122,300 as dividends.